OpenStudy (anonymous):

solve using the multiplication property or identity 6/7y= 2/3

OpenStudy (anonymous):

multiply by 7/6 y = 14/18 = 7/9

OpenStudy (anonymous):

\[\frac{ 6 }{ 7 }y=\frac{ 2 }{ 3 }\]is it your question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can divide both of side with \[\frac{ 6 }{ 7 }\]

OpenStudy (anonymous):

dont u multyply them

OpenStudy (anonymous):

so, the equation will be\[\frac{ \frac{ 6 }{ 7 } }{ \frac{ 6 }{ 7 } }y=\frac{ \frac{ 2 }{ 3 } }{ \frac{ 6 }{ 7 } }\]\[y=\frac{ 2 }{ 3 }\times \frac{ 7 }{ 6 }\]now, you can solve by yourself

OpenStudy (anonymous):

you actually have to multiply both sides by 6/7

OpenStudy (anonymous):

its that, multiply both the sides by 7/6

OpenStudy (anonymous):

not multiply, but divide. if you multiply both of side with 6/7, the result will be \[(\frac{ 6 }{ 7 })^{2}y=\frac{ 2 }{ 3 }\times \frac{ 6 }{ 7 }\]it will not solve your problem then

OpenStudy (anonymous):

it actually does they never told us nothing about dividing in school

OpenStudy (anonymous):

actually it is about how you can eliminate 6/7 from that equation, so only y remain (in the left side).

OpenStudy (anonymous):

if they taught to multiply, you need to multiply 6/7 by its reciprocal, i.e. 7/6

OpenStudy (anonymous):